What are four decimal fractions that are smaller than 0.5
What are four vulgar fractions that are greater than 1/2 but smaller than 3/4

Answer by richwmiller(17219)

What are four decimal fractions that are smaller than 0.5
.1
.2
.3
.4
What are four vulgar fractions that are greater than 1/2 but smaller than 3/4
12/24=1/2
18/24=3/4
13/24
14/24 = 7/12
15/24
16/24=2/3